Circular No. 6934

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998cl IN MARKARIAN 261 Michael Schwartz, Cottage Grove, OR, reports the discovery of an ap-

parent supernova (mag ∼ 17) located at α = 13^{h}29^{m}59^{s}.8, δ = +75^{o}34′09′′.0 (equinox 2000.0), which is ∼ 5′′ west and 1′′ north of Markarian 261 (whose nucleus is ill-defined) on unfiltered CCD frames taken with the 0.35-m Tenagra I automated supernova patrol telescope on June 3.28 UT. The object is also present with unchanged brightness on a frame taken with the same equipment on June 6.26. The object does not appear on an unfiltered CCD image taken with the same equipment on 1997 July 7.33, and the star is not present on second Palomar Sky Survey plates.

XTE J1748–288 T. Strohmayer and F. E. Marshall, Goddard Space Flight Center, on

behalf of the RXTE Science Operations Center, report an improved position for this x-ray transient (IAUC 6932) of α = 17^{h}48^{m}08^{s}, δ = −28^{o}29′.2 (equinox 2000.0) with an estimated uncertainty of 1′ based on scans across the source with the Proportional Counter Array on the Rossi X-ray Timing Explorer on June 5.13 UT. The observed 2–10-keV flux of ∼ 600 mCrab indicates a continued brightening of the source.

R. M. Hjellming and M. P. Rupen, National Radio Astronomy Observatory (NRAO); and A. J. Mioduszewski, Joint Institute for VLBI in Europe and NRAO, report: “Very Large Array observations on June 7.40 UT of the field of XTE J1748–288 show the presence of an unresolved radio source with flux densities of 50 ± 3 and 28 ± 2 mJy at 1.46 and 4.86 GHz, respectively. Its position is α = 17^{h}48^{m}05^{s}.06, δ = −28^{o}28′25′′.8 (equinox 2000.0; uncertainty 0′′.6), which is 1′ from the center of the 1′ uncertainty region for XTE J1748–288 (see above). It is a candidate for the radio counterpart of the x-ray transient because it is much stronger than the usual background radio source, but variability still needs to be established.”

COMET C/1998 J1 (SOHO) J. Crovisier, D. Bockelée-Morvan, P. Colom, and E. Gérard, Observa-

toire de Paris-Meudon; and N. Biver, Institute for Astronomy, University of Hawaii, report observations of OH 18-cm lines with the Nançay radio telescope around the time of the outburst of the comet (cf. IAUC 6926). The OH production rates were as follows, in units of 10^{29} molecules/s: June 1.6 UT, 3; 2.6, 5; 3.6, 4; 4.6, 3.

1998 June 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green